# Topology of Discrete Quantum Feedback Control

### v2

@@ Acknowledgements.
M.N. acknowledges helpful discussions with Yuto Ashida, Kohei Kawabata, Masatoshi Sato, Kenji Shimomura, and Ken Shiozaki during the workshop YITP-T-24-03 on “Recent Developments and Challenges in Topological Phases” held at the Yukawa Institute for Theoretical Physics. This work was supported by KAKENHI Grant No. JP20K14383, No. JP22H01152, and No. JP24K16989 from the Japan Society for the Promotion of Science. We gratefully acknowledge the support from the CREST program “Quantum Frontiers” (Grant No. JPMJCR23I1) by the Japan Science and Technology Agency.
@@ Appendix A Bloch theory of quantum channels
The Bloch theory lies at the heart of topological band theory [1]. However, how to generalize it to feedback-controlled quantum systems is nontrivial since feedback control involves not only unitary time evolution but also measurement processes. Here, we develop a Bloch theory of quantum channels described by CPTP maps. The following formalism is not limited to discrete quantum feedback control but can generally be applied to any translationally invariant quantum channels in single-particle systems.
@@ A.1 CPTP map as a non-Hermitian operator
We begin with representing a CPTP map in terms of a non-Hermitian matrix. We consider a single-particle quantum system on a lattice (see Sec. III.1 for details). We first vectorize the density matrix
@@
Namely, we have the matrix representation of $\mathcal{E}$ as
@@ A.2 Momentum-space representation
Here we consider a translationally invariant quantum channel satisfying
@@
for arbitrary $\hat{\rho}$ with $\hat{T}_{\lambda}\ (\lambda=1,\cdots,d)$ being a translation operator:
@@
In the context of open quantum systems, this symmetry can be regarded as weak symmetry of the CPTP map (see Sec. V.1) [92, 93]. In fact, since the operator $\tilde{\mathcal{E}}$ commutes with the tensor-product translation operator $\hat{T}_{\lambda}\otimes\hat{T}_{\lambda}^{*}$ appearing in Eq. (202), the operator $\tilde{\mathcal{E}}$ can be block-diagonalized with respect to eigenvalues of $\hat{T}_{\lambda}\otimes\hat{T}_{\lambda}^{*}$. Eigenvectors of $\hat{T}_{\lambda}\otimes\hat{T}_{\lambda}^{*}$ are constructed from the eigenstates
@@
Thus, eigenvectors of $\tilde{\mathcal{E}}$ are classified according to the eigenvalues of $\hat{T}_{\lambda}\otimes\hat{T}_{\lambda}^{*}$. Since the vector |$\ket{\vec{k},a}\otimes\ket{-\vec{k}^{\prime},b}$ corresponds to an operator |$\ket{\vec{k},a}\bra{\vec{k}^{\prime},b}$ | through the Choi-Jamiołkowski isomorphism (see Eqs. (193) and (194)), $\vec{k}-\vec{k}^{\prime}$ in Eq. (204) is physically interpreted as the difference between crystal momenta of the ket and bra degrees of freedom of the density matrix.
@@
For later convenience, we assume that components $\mu_{\lambda}$ and $\nu_{\lambda}$ in $\vec{\mu}=(\mu_{1},\cdots,\mu_{d})$ and $\vec{\nu}=(\nu_{1},\cdots,\nu_{d})$ take a value from $-L/2,-L/2+1,\cdots,L/2-1$ ($-(L-1)/2,-(L-1)/2+1,\cdots,(L-1)/2$) for even (odd) $L$ without loss of generality by using the periodic boundary condition (20). The translational symmetry (202) exists if
@@
Here, we note that we consider a single-particle system although Eq. (210) is written like the second-quantized form. We also note that the operator $\tilde{c}_{\vec{i},a,c,\vec{\mu}}^{\dagger}$ has two additional indices $c$ and $\vec{\mu}$ compared with the original basis |$\ket{\vec{i},a}$ of the Hilbert space since $\tilde{\mathcal{E}}$ acts on the doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$. Physically, this corresponds to the fact that the density matrix has bra and ket degrees of freedom.
@@
The $N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}\times N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}$ matrix $X(\vec{k})$ is a counterpart of the Bloch Hamiltonian in the band theory, and therefore we shall call it the Bloch matrix.
@@ A.3 Mode expansion
Let us express the density matrix $\hat{\rho}_{\mathrm{f}}=\mathcal{E}(\hat{\rho}_{\mathrm{i}})$ after feedback control in terms of eigenvalues and eigenvectors of the Bloch matrix $X(\vec{k})$. Let $v_{n}^{L}(\vec{k})$ ($v_{n}^{R}(\vec{k})$) be a left (right) eigenvector of $X(\vec{k})$ with eigenvalue $\xi_{n}(\vec{k})$. We assume that the Bloch matrix $X(\vec{k})$ is diagonalizable and that the eigenvectors satisfy the biorthogonal condition $(v_{m}^{L}(\vec{k}),v_{n}^{R}(\vec{k}))=\delta_{m,n}$ (here $(u,v)$ denotes the inner product of vectors $u$ and $v$) and the completeness relation $\sum_{n}v_{n}^{R}(\vec{k})(v_{n}^{L}(\vec{k}))^{\dagger}=1$. Then, the Bloch matrix is written as
@@
Substituting this into Eq. (211), we have
@@ A.4 Fundamental constraints on the Bloch matrix
Since a quantum channel must satisfy the complete positivity and the trace preservation, the Bloch matrix $X(\vec{k})$ cannot be a general non-Hermitian matrix. We note that a superoperator satisfies the complete positivity if it has the Kraus representation (5) (see e.g., Ref. [80]). The Kraus representation implies the Hermiticity preservation, which leads to an antiunitary symmetry (62) of a quantum channel (see Sec. V.2). Using Eq. (210), the right-hand side of the symmetry condition (62) is written as
@@
Next, we consider the condition for the trace preservation. The superoperator preserves the trace of the density matrix if
@@ Appendix B Feedback control with projective measurement
Here we consider an example of quantum feedback control with a projective measurement, where the measurement operators are given by $\hat{M}_{m}=\ket{\vec{j},a}\bra{\vec{j},a}$ | and a measurement outcome is denoted by $m=(\vec{j},a)$. In this case, the matrix elements of the Kraus operator $\hat{K}_{m}=\hat{U}_{m}\hat{M}_{m}$ are written as
@@
In particular, if a system has no internal degrees of freedom (i.e., $D_{\mathrm{loc}}=1$), the analysis is further simplified. In this case, we can omit the indices for internal degrees of freedom and the matrix elements of the Bloch matrix are written as
@@
Since the Bloch matrix has nonzero matrix elements only in one columun, its eigenvalues $\xi_{\vec{\mu}}(\vec{k})$ are given by its diagonal entries as
@@
The Bloch matrix for the $D_{\mathrm{loc}}=1$ case is diagonalizable if $X_{\vec{0},\vec{0}}(\vec{k})\neq 0$. If $X_{\vec{0},\vec{0}}(\vec{k})=0$, it cannot be diagonalized except for the trivial case with $X(\vec{k})=0$. Thus, the point at which $X_{\vec{0},\vec{0}}(\vec{k})=\xi_{\vec{0}}(\vec{k})=0$ corresponds to an exceptional (non-diagonalizable) point [68]. Specifically, if $X_{\vec{0},\vec{0}}(\vec{k})\neq 0$, the left eigenvector $v_{\vec{\nu}}^{L}(\vec{k})$ and the right eigenvector $v_{\vec{\nu}}^{R}(\vec{k})$ corresponding to the eigenvalue $\xi_{\vec{\nu}}(\vec{k})$ are given by
@@
are determined from the probability distribution $p_{\vec{j}}=\mathrm{Tr}[\hat{M}_{\vec{j}}^{\dagger}\hat{M}_{\vec{j}}\hat{\rho}_{\mathrm{i}}]$ of measurement outcomes. The probability distribution of the position after the feedback control is given by
@@ Appendix C Eigenvalues of a quantum channel for projective measurement
Here we show that eigenvalues of a quantum channel for projective measurement are either zero or one. We consider a quantum channel
@@
for an arbitrary operator $\hat{A}$. Thus, eigenvalues of $\mathcal{E}_{\mathrm{proj}}$ are real and nonnegative. These eigenvalues must be either zero or one since the measurement channel satisfies the idempotency condition:
@@ Appendix D Correspondence between symmetry of a time-evolution operator and that of its generator
Suppose that a time-evolution operator has a generator. In order to investigate the symmetry of a time-evolution operator, it is helpful to consider the symmetry of its generator. Here we summarize the correspondence between symmetry of a time-evolution operator and that of its generator to compare the former with symmetry of quantum channels discussed in Sec. V.3.
In closed quantum systems, a time evolution operator $\hat{U}=e^{-i\hat{H}t}$ is generated by a Hermitian Hamiltonian $H$. The Altland-Zirnbauer symmetries of a Hermitian Hamiltonian are written as
@@
$-\hat{\Gamma}H\hat{\Gamma}^{-1}=$ $H,\ \hat{\Gamma}^{2}=\hat{I},$ (251c)
@@
In Markovian open quantum systems, the generator (Lindbladian) $\mathcal{L}$ of a CPTP map $\mathcal{E}=e^{\mathcal{L}t}$ for time evolution features the Gorini-Kossakowski-Sudarshan-Lindblad form [99]. The matrix representation of a CPTP map through the Choi-Jamiołkowski isomorphism (see Sec. A.1) is denoted by ~$\tilde{\mathcal{E}}=e^{\tilde{\mathcal{L}}t}$. Here, we note that a Lindbladian cannot be a general non-Hermitian matrix since the real part of its eigenvalues must be non-positive. Because of this constraint, only the following set of Bernard-LeClair symmetries are possible for a Lindbladian [65] (see also Refs. [66, 67]):
@@ Appendix E Symmetry-preserving boundary condition for a $\mathbb{Z}_{2}$ demon
In this appendix, we describe how to implement a symmetry-preserving boundary condition for the $\mathbb{Z}_{2}$ demon in Sec. VII.4. The transition probability amplitudes for a feedback control in class AII must satisfy Eqs. (111a)-(111c). Here we perform a feedback control near the boundary of the system to preserve these symmetry conditions. We consider a 1D lattice with $L$ unit cells under the OBC, and the site $j=1$ ($j=L$) denotes the left (right) boundary. The feedback control under the PBC satisfies the conditions (111a)-(111c). However, if we perform the same feedback under the OBC, the reflection of a particle at the boundary violates these conditions. A feedback protocol that preserves the symmetry at the boundary is given as follows. Let $m_{1}=(j_{1},a_{1})$ be a measurement outcome of the first measurement. If $j_{1}=1$ or $j_{1}=L$, we do nothing (i.e., the feedback Hamiltonian is set to zero: $\hat{H}_{1,a_{1}}=\hat{H}_{L,a_{1}}=0$). The transition amplitudes from the boundary sites therefore vanish. If $j_{1}\neq 1,L$, we perform the feedback control with the same Hamiltonian $\hat{H}_{j_{1},a_{1}}$ as that for the PBC case described in Sec. VII.4. Then, we perform the second projective measurement and obtain an outcome $m_{2}=(j_{2},a_{2})$. If $j_{2}=2,\cdots,L-1$, we perform the same additional feedback control as the PBC case shown in Eq. (155). If $j_{2}=1$ or $j_{2}=L$, we perform a feedback operation with the unitary operator
@@
The existence of a solution of Eqs. (258) and (259) depends on the detail of the transition probability (260). For example, if
@@
A solution of Eq. (259) can be obtained similarly. For the parameters employed in the numerical simulation in Fig. 18, the conditions (261) and (262) are met, and therefore we set the phase factors as Eqs. (263)-(265) in Fig. 18.
It is worthwhile to note that conditions (258) and (259) may not have a solution. In that case, the symmetry of a quantum channel should be violated under the OBC and the symmetry-protected non-Hermitian skin effect is not ensured. However, if the symmetry is broken only near the boundary, the transport of a particle towards the boundary due to the skin effect is partially maintained for a certain time before a particle approaches the boundary, since the dynamics in the bulk is unaltered [33].
Figure 23: Schematic illustration of topological feedback control that shows the non-Hermitian skin effect but with no boundary-localized steady state. A two-level atom with internal states $g,e$ on a 1D lattice is considered. The excited state $e$ spontaneously decays into the ground state $g$ with rate $\gamma$. If a projective measurement finds the atom in the excited state, a potential wall with height $V$ is introduced on the left site to the particle.
Figure 24: Eigenspectrum of a quantum channel in Eq. (268). The parameters are set to be $J_{e}=1,J_{g}=0.2,V=10,\gamma=2$, and $\tau=1$. Blue curves and blue dots at the origin constitue the eigenspectrum under the PBC for infinite system size, and orange dots show the eigenspectrum under the OBC for the system size $L=20$.
@@
The models of topological feedback control presented in Sec. IV, VII.1, VII.2, VII.3, VII.4, and IX show the non-Hermitian skin effects and boundary-localized steady states due to nontrivial point-gap topology of quantum channels. While the former is widely observed in non-Hermitian topological phases [33, 72, 73, 88, 38, 74], the latter does not exist in conventional non-Hermitian systems and thus is unique to topological quantum channels which preserve the positivity and trace of a density matrix (see Sec. VIII.1). Here, to distinguish these two topological phenomena, we present a model that shows the non-Hermitian skin effect but with no boundary-localized steady states. We consider a two-level atom on a 1D lattice with $L$ sites. The internal states of the atom consist of a ground state ($g$) and an excited state ($e$). We perform a projective measurement of the position and the internal state of the atom described by the measurement operator $\hat{M}_{j,a}=\ket{j,a}\bra{j,a}$ |, where |$\ket{j,a}$ denotes a quantum state of the atom at site $j=1,\cdots,L$ and internal state $a=g,e$. Then, depending on the measurement outcome $m=(j,a)$, we let the atom undergo the time evolution under the Lindblad equation
@@
is the feedback Hamiltonian with hopping amplitudes $J_{e},J_{g}\in\mathbb{R}$ and potential height $V>0$, and $\hat{L}_{i}=\sqrt{\gamma}\ket{i,g}\bra{i,e}$ | is a Lindblad operator that describes spontaneous decay with rate $\gamma>0$ from the excited state to the ground state. That is, we impose a potential wall only if the internal state of the atom is detected to be the excited state. See Fig. 23 for a schematic illustration. The quantum channel for the feedback control is thus given by (see also Sec. IX)
@@
where $\tau$ is the duration of the feedback control.
Figure 24 shows the eigenspectrum of the quantum channel (268). Here, a salient feature is that the band in the PBC spectrum that involves the steady state (the unit eigenvalue) does not form a loop and thus has a vanishing winding number. Physically, this is because the atom is in the internal ground state in the steady state and thus is not localized near the edges of the system. Meanwhile, the band located away from the steady state exhibits a loop structure. This leads to the non-Hermitian skin effect of an atom in the internal excited state, which has a finite lifetime due to spontaneous decay. Thus, the non-Hermitian skin effect in topological feedback control is not always accompanied by a boundary-localized steady state.

### v3

@@ Acknowledgements.
M.N. acknowledges helpful discussions with Yuto Ashida, Kohei Kawabata, Masatoshi Sato, Kenji Shimomura, and Ken Shiozaki during the workshop YITP-T-24-03 on “Recent Developments and Challenges in Topological Phases” held at the Yukawa Institute for Theoretical Physics. This work was supported by KAKENHI Grants No. JP20K14383, No. JP22H01152, and No. JP24K16989 from the Japan Society for the Promotion of Science. We gratefully acknowledge support from the CREST program “Quantum Frontiers” (Grant No. JPMJCR23I1) by the Japan Science and Technology Agency.
@@ Appendix A Bloch theory of quantum channels
The Bloch theory lies at the heart of topological band theory [1]. However, how to generalize this theory to feedback-controlled quantum systems is nontrivial since feedback control involves not only unitary time evolution but also measurement processes. Here, we develop a Bloch theory of quantum channels described by CPTP maps. The following formalism is not limited to discrete quantum feedback control but can generally be applied to any translationally invariant quantum channels in single-particle systems.
@@ A.1 CPTP map as a non-Hermitian operator
We begin by representing a CPTP map in terms of a non-Hermitian matrix. We consider a single-particle quantum system on a lattice (see Sec. III.1 for details). We first vectorize the density matrix
@@
Thus, the matrix representation of $\mathcal{E}$ is
@@ A.2 Momentum-space representation
Here, we consider a translationally invariant quantum channel satisfying
@@
for arbitrary $\hat{\rho}$, with $\hat{T}_{\lambda}\ (\lambda=1,\cdots,d)$ being a translation operator:
@@
In the context of open quantum systems, this symmetry can be regarded as weak symmetry of the CPTP map (see Sec. V.1) [92, 93]. In fact, since the operator $\tilde{\mathcal{E}}$ commutes with the tensor-product translation operator $\hat{T}_{\lambda}\otimes\hat{T}_{\lambda}^{*}$ appearing in Eq. (202), the operator $\tilde{\mathcal{E}}$ can be block diagonalized with respect to eigenvalues of $\hat{T}_{\lambda}\otimes\hat{T}_{\lambda}^{*}$. Eigenvectors of $\hat{T}_{\lambda}\otimes\hat{T}_{\lambda}^{*}$ are constructed from the eigenstates
@@
Thus, eigenvectors of $\tilde{\mathcal{E}}$ are classified according to the eigenvalues of $\hat{T}_{\lambda}\otimes\hat{T}_{\lambda}^{*}$. Since the vector |$\ket{\vec{k},a}\otimes\ket{-\vec{k}^{\prime},b}$ corresponds to an operator |$\ket{\vec{k},a}\bra{\vec{k}^{\prime},b}$ | through the Choi-Jamiołkowski isomorphism [see Eqs. (193) and (194)], $\vec{k}-\vec{k}^{\prime}$ in Eq. (204) is physically interpreted as the difference between crystal momenta of the ket and bra degrees of freedom of the density matrix.
@@
For later convenience, we assume that components $\mu_{\lambda}$ and $\nu_{\lambda}$ in $\vec{\mu}=(\mu_{1},\cdots,\mu_{d})$ and $\vec{\nu}=(\nu_{1},\cdots,\nu_{d})$ take a value from $-L/2,-L/2+1,\cdots,L/2-1$ [$-(L-1)/2,-(L-1)/2+1,\cdots,(L-1)/2$] for even (odd) $L$ without loss of generality by using the periodic boundary condition (20). The translational symmetry (202) exists if
@@
Here, we consider a single-particle system, although Eq. (210) is written like the second-quantized form. We also note that the operator $\tilde{c}_{\vec{i},a,c,\vec{\mu}}^{\dagger}$ has two additional indices $c$ and $\vec{\mu}$ compared with the original basis |$\ket{\vec{i},a}$ of the Hilbert space since $\tilde{\mathcal{E}}$ acts on the doubled Hilbert space $\mathcal{H}\otimes\mathcal{H}$. Physically, this point corresponds to the fact that the density matrix has bra and ket degrees of freedom.
@@
The $N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}\times N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}$ matrix $X(\vec{k})$ is a counterpart of the Bloch Hamiltonian in band theory; therefore, we call it the Bloch matrix.
@@ A.3 Mode expansion
Let us express the density matrix $\hat{\rho}_{\mathrm{f}}=\mathcal{E}(\hat{\rho}_{\mathrm{i}})$ after feedback control in terms of eigenvalues and eigenvectors of the Bloch matrix $X(\vec{k})$. Let $v_{n}^{L}(\vec{k})$ [$v_{n}^{R}(\vec{k})$] be a left (right) eigenvector of $X(\vec{k})$ with eigenvalue $\xi_{n}(\vec{k})$. We assume that the Bloch matrix $X(\vec{k})$ is diagonalizable and that the eigenvectors satisfy the biorthogonal condition $(v_{m}^{L}(\vec{k}),v_{n}^{R}(\vec{k}))=\delta_{m,n}$ [here, $(u,v)$ denotes the inner product of vectors $u$ and $v$] and the completeness relation $\sum_{n}v_{n}^{R}(\vec{k})(v_{n}^{L}(\vec{k}))^{\dagger}=1$. Then, the Bloch matrix is written as
@@
Substituting this result into Eq. (211), we have
@@ A.4 Fundamental constraints on the Bloch matrix
Since a quantum channel must satisfy the complete positivity and the trace preservation, the Bloch matrix $X(\vec{k})$ cannot be a general non-Hermitian matrix. We note that a superoperator satisfies the complete positivity if it has the Kraus representation (5) (see e.g., Ref. [80]). The Kraus representation implies Hermiticity preservation, which leads to an antiunitary symmetry (62) of a quantum channel (see Sec. V.2). Using Eq. (210), the right-hand side of the symmetry condition (62) is written as
@@
Next, we consider the condition for trace preservation. The superoperator preserves the trace of the density matrix if
@@ Appendix B Feedback control with projective measurement
Here we consider quantum feedback control with a projective measurement, where the measurement operators are given by $\hat{M}_{m}=\ket{\vec{j},a}\bra{\vec{j},a}$ | and a measurement outcome is denoted by $m=(\vec{j},a)$. In this case, the matrix elements of the Kraus operator $\hat{K}_{m}=\hat{U}_{m}\hat{M}_{m}$ are written as
@@
In particular, if a system has no internal degrees of freedom (i.e., $D_{\mathrm{loc}}=1$), the analysis is further simplified. In this case, we can omit the indices for internal degrees of freedom, and the matrix elements of the Bloch matrix are written as
@@
Since the Bloch matrix has nonzero matrix elements only in one column, its eigenvalues $\xi_{\vec{\mu}}(\vec{k})$ are given by its diagonal entries as
@@
The Bloch matrix for the $D_{\mathrm{loc}}=1$ case is diagonalizable if $X_{\vec{0},\vec{0}}(\vec{k})\neq 0$. If $X_{\vec{0},\vec{0}}(\vec{k})=0$, it cannot be diagonalized except for the trivial case with $X(\vec{k})=0$. Thus, the point at which $X_{\vec{0},\vec{0}}(\vec{k})=\xi_{\vec{0}}(\vec{k})=0$ corresponds to an exceptional (nondiagonalizable) point [68]. Specifically, if $X_{\vec{0},\vec{0}}(\vec{k})\neq 0$, the left eigenvector $v_{\vec{\nu}}^{L}(\vec{k})$ and the right eigenvector $v_{\vec{\nu}}^{R}(\vec{k})$ corresponding to the eigenvalue $\xi_{\vec{\nu}}(\vec{k})$ are given by
@@
are determined by the probability distribution $p_{\vec{j}}=\mathrm{Tr}[\hat{M}_{\vec{j}}^{\dagger}\hat{M}_{\vec{j}}\hat{\rho}_{\mathrm{i}}]$ of measurement outcomes. The probability distribution of the position after feedback control is given by
@@ Appendix C Eigenvalues of a quantum channel for projective measurement
Here, we show that eigenvalues of a quantum channel for a projective measurement are either zero or one. We consider a quantum channel
@@
for an arbitrary operator $\hat{A}$. Thus, eigenvalues of $\mathcal{E}_{\mathrm{proj}}$ are real and non-negative. These eigenvalues must be either zero or one since the measurement channel satisfies the idempotency condition:
@@ Appendix D Correspondence between symmetry of a time-evolution operator and that of its generator
Suppose that a time-evolution operator has a generator. To investigate the symmetry of a time-evolution operator, it is helpful to consider the symmetry of its generator. Here, we summarize the correspondence between symmetry of a time-evolution operator and that of its generator to compare the former with the symmetry of quantum channels discussed in Sec. V.3.
In closed quantum systems, a time-evolution operator $\hat{U}=e^{-i\hat{H}t}$ is generated by a Hermitian Hamiltonian $\hat{H}$. The Altland-Zirnbauer symmetries of a Hermitian Hamiltonian are written as
@@
$-\hat{\Gamma}\hat{H}\hat{\Gamma}^{-1}=$ $\hat{H},\ \hat{\Gamma}^{2}=\hat{I},$ (251c)
@@
In Markovian open quantum systems, the generator (Lindbladian) $\mathcal{L}$ of a CPTP map $\mathcal{E}=e^{\mathcal{L}t}$ for time evolution features the Gorini-Kossakowski-Sudarshan-Lindblad form [99]. The matrix representation of a CPTP map through the Choi-Jamiołkowski isomorphism (see Appendix A.1) is denoted by ~$\tilde{\mathcal{E}}=e^{\tilde{\mathcal{L}}t}$. Here, we note that a Lindbladian cannot be a general non-Hermitian matrix since the real part of its eigenvalues must be nonpositive. Because of this constraint, only the following set of Bernard-LeClair symmetries are possible for a Lindbladian [65] (see also Refs. [66, 67]):
@@ Appendix E Symmetry-preserving boundary condition for a $\mathbb{Z}_{2}$ demon
In this appendix, we describe how to implement a symmetry-preserving boundary condition for the $\mathbb{Z}_{2}$ demon in Sec. VII.4. The transition probability amplitudes for a feedback control in class AII must satisfy Eqs. (111a)-(111c). Here we perform a feedback control near the boundary of the system to preserve these symmetry conditions. We consider a 1D lattice with $L$ unit cells under the OBC, and the site $j=1$ ($j=L$) denotes the left (right) boundary. Feedback control under the PBC satisfies the conditions (111a)-(111c). However, if we perform the same feedback under the OBC, the reflection of a particle at the boundary violates these conditions. A feedback protocol that preserves the symmetry at the boundary is given as follows. Let $m_{1}=(j_{1},a_{1})$ be a measurement outcome of the first measurement. If $j_{1}=1$ or $j_{1}=L$, we do nothing (i.e., the feedback Hamiltonian is set to zero: $\hat{H}_{1,a_{1}}=\hat{H}_{L,a_{1}}=0$). The transition amplitudes from the boundary sites therefore vanish. If $j_{1}\neq 1,L$, we perform feedback control with the same Hamiltonian $\hat{H}_{j_{1},a_{1}}$ as that for the PBC case described in Sec. VII.4. Next, we perform the second projective measurement and obtain an outcome $m_{2}=(j_{2},a_{2})$. If $j_{2}=2,\cdots,L-1$, we perform the same additional feedback control as the PBC case shown in Eq. (155). If $j_{2}=1$ or $j_{2}=L$, we perform a feedback operation with the unitary operator
@@
The existence of a solution of Eqs. (258) and (259) depends on the details of the transition probability (260). For example, if
@@
A solution of Eq. (259) can be obtained similarly. For the parameters employed in the numerical simulation in Fig. 18, the conditions (261) and (262) are met, and therefore, we set the phase factors as Eqs. (263)-(265) in Fig. 18.
It is worthwhile to note that conditions (258) and (259) may not have a solution. In that case, the symmetry of a quantum channel should be broken under the OBC, and the symmetry-protected non-Hermitian skin effect is not ensured. However, if the symmetry is broken only near the boundary, the transport of a particle towards the boundary due to the skin effect is partially maintained for a certain time before a particle approaches the boundary since the dynamics in the bulk is unaltered [33].
Figure 23: Schematic illustration of topological feedback control that shows the non-Hermitian skin effect but with no boundary-localized steady state. A two-level atom with internal states $g$ and $e$ on a 1D lattice is considered. The excited state $e$ spontaneously decays into the ground state $g$ with rate $\gamma$. If a projective measurement finds the atom in the excited state, a potential wall with height $V$ is introduced on the left side of the particle.
Figure 24: Eigenspectrum of a quantum channel in Eq. (268). The parameters are set to $J_{e}=1,J_{g}=0.2,V=10,\gamma=2$, and $\tau=1$. Blue curves and blue dots at the origin constitute the eigenspectrum under the PBC for infinite system size, and orange dots show the eigenspectrum under the OBC for the system size $L=20$.
@@
The models of topological feedback control presented in Secs. IV, VII.1-VII.4, and IX show the non-Hermitian skin effects and boundary-localized steady states due to nontrivial point-gap topology of quantum channels. While the former is widely observed in non-Hermitian topological phases [33, 72, 73, 88, 38, 74], the latter does not exist in conventional non-Hermitian systems and thus is unique to topological quantum channels that preserve the positivity and trace of a density matrix (see Sec. VIII.1). Here, to distinguish these two topological phenomena, we present a model that shows the non-Hermitian skin effect but with no boundary-localized steady states. We consider a two-level atom on a 1D lattice with $L$ sites. The internal states of the atom consist of a ground state ($g$) and an excited state ($e$). We perform a projective measurement of the position and the internal state of the atom described by the measurement operator $\hat{M}_{j,a}=\ket{j,a}\bra{j,a}$ |, where |$\ket{j,a}$ denotes a quantum state of the atom at site $j=1,\cdots,L$ and internal state $a=g,e$. Next, depending on the measurement outcome $m=(j,a)$, we let the atom undergo time evolution under the Lindblad equation
@@
is the feedback Hamiltonian with hopping amplitudes $J_{e},J_{g}\in\mathbb{R}$ and potential height $V>0$, and $\hat{L}_{i}=\sqrt{\gamma}\ket{i,g}\bra{i,e}$ | is a Lindblad operator that describes spontaneous decay with rate $\gamma>0$ from the excited state to the ground state. In other words, we impose a potential wall only if the internal state of the atom is detected to be the excited state. See Fig. 23 for a schematic illustration. The quantum channel for feedback control is thus given by (see also Sec. IX)
@@
where $\tau$ is the duration of feedback control.
Figure 24 shows the eigenspectrum of the quantum channel (268). Here, a salient feature is that the band in the PBC spectrum that involves the steady state (the unit eigenvalue) does not form a loop and thus has a vanishing winding number. Physically, this feature occurs because the atom is in the internal ground state in the steady state and thus is not localized near the edges of the system. Meanwhile, the band located away from the steady state exhibits a loop structure. This topological structure leads to the non-Hermitian skin effect of an atom in the internal excited state, which has a finite lifetime due to spontaneous decay. Thus, the non-Hermitian skin effect in topological feedback control is not always accompanied by a boundary-localized steady state.